A bank loaned out $12,000, part of it at the rate of 8% per year and the rest at the rate of 18% per year. If the interest recieved totaled $1,000, how much was loaned at 8%?

Answer by checkley71(8403)

.08x+.18(10000-x)=1000
.08x+1800-.18x=1000
-.1x=1000-1800
-.1x=-800
x=-800/-.1
------------------------------------------------------
x=8000 amount loaned @ 8%.
-----------------------------------------------------
10000-8000=2000 loaned @ 18%.
proof
8000*.08+2000*.18=1000
640+360=1000
1000=1000